# Correlation functions of harmonic lattices in d-dimensional space

### v2

@@ 1 Introduction
In this manuscript, we show that by considering the thermodynamic limit, the lattice scale $L\to\infty$ and replacing the sum over all site positions $(k_{1},k_{2},\ldots,k_{d})$ within $U$ appearing in the computation of $Q$ and $P$ with multiple integrals $\int\int\dots\int dk_{1}dk_{2}\dots dk_{d}$ and calculating them analytically, the results for any spatial dimension expressed by the C-type Lauricella’s hypergeometric functions, $F_{C}$, irrespective of the boundary conditions for the dynamical variables $q_{\vec{r}}$s and their conjugate momenta $p_{\vec{r}}$s. We expect that by utilizing these expressions along with appropriate numerical computation libraries, it makes easier to create programs that rapidly compute various information quantities arising from entanglement.
@@ 4 Connection between correlators of different boundary conditions
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim(e^{-N_{+}}\prod_{\alpha=1}^{d_{1}}n_{\alpha}^{n_{\alpha}+1/2}e^{-N_{+}})^{-1}\sim(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{\lambda_{\alpha}N_{+}})^{-1}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (29)

### v3

@@ 1 Introduction
In this manuscript, we show that by considering the thermodynamic limit, the lattice scale $L\to\infty$ and replacing the sum over all site positions $(k_{1},k_{2},\ldots,k_{d})$ within $U$ appearing in the computation of $Q$ and $P$ with multiple integrals $\int\int\dots\int dk_{1}dk_{2}\dots dk_{d}$ and calculating them analytically, the results for any spatial dimension are expressed by the C-type Lauricella’s hypergeometric functions, $F_{C}$, irrespective of the boundary conditions for the dynamical variables $q_{\vec{r}}$s and their conjugate momenta $p_{\vec{r}}$s. We expect that by utilizing these expressions along with appropriate numerical computation libraries, it makes easier to create programs that rapidly compute various information quantities arising from entanglement.
@@ 4 Connection between correlators of different boundary conditions
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim(e^{-N_{+}}\prod_{\alpha=1}^{d_{1}}n_{\alpha}^{n_{\alpha}+1/2})^{-1}\sim(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{\lambda_{\alpha}N_{+}})^{-1}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (29)